DIRECTIONS: Solve the following.

1. Determine the number of atoms in 1.75 mol Al

2. Calculate the number of formula units in 25.0 mol NaNO3

3. Given 6.25 mol H O2 determine the number of molecules.



1
Expert's answer
2021-06-01T02:44:59-0400

Mole="\frac{Mass}{Molar mass}"


Molarity= "\frac{Mass of solute}{ volume in liters}"


Molality= "\frac{moles of solute} {volume of solvent in kg}"

These are the common formulas related to moles that are used in chemical calculations. 

1) When aluminum has 1.75 moles, we need to calculate the number of atoms in it. For calculating the number of moles, the value is multiplied by the Avogadro's number.

One mole=6.022×1023

Number of atoms="\frac{1.75mol}{1mol}" ×6.022×1023atoms

Number of atoms =1.05×1024atoms



2) In sodium nitrate, 25.0 moles are present. To calculate the number of formula units in it, we need to multiply the moles value by Avogadro's number.

One mole=6.022×1023

Number of formula units ="\frac{25.0mol}{1mol}" ×6.022×1023formula units


Number of formula units =1.51×1025formula units


3) In sodium nitrate, 25.0 moles are present. To calculate the number of formula units in it, we need to multiply the moles value by Avogadro's number.

One mole =6.022×1023

Number of molecules ="\frac{6.25mol}{1mol}" ×6.022×1023molecules

Number of molecules =3.76×1024molecules